OpenStudy (anonymous):

Solve for x. 2x + 16 = 3(x - 9) A) -43 B) -25 C) 25 D) 43

OpenStudy (anonymous):

Distribute first

OpenStudy (anonymous):

2x + 16 = 3(x - 9) so, 2x + 16 = 3x - 27

OpenStudy (anonymous):

Now combine like terms

OpenStudy (anonymous):

3x - 2x = 16 + 27

OpenStudy (anonymous):

x = 4. Your answer is D